DYNAMICS OF THRESHOLD PHOTO-ION PAIR PRODUCTION IN SMALL MOLECULES

\underline{JOHN W. HEPBURN}, QICHI HU, University of British Columbia, Vancouver, BC, Canada, V6T 1Z4.

It has already been demonstrated that the method of threshold ion-pair production spectroscopy (TIPPS) can be used to probe the energetics of molecular dissociation with spectroscopic accuracy. Since TIPPS produces a spectrum which measures energy differences between the starting neutral molecule states and the states of the ion-pair dissociation products, it can also be used to determine the spectroscopy of the ion fragments, and the dynamics of the ion-pair dissociation. The results of TIPPS experiments on the following dissociations will be used to illustrate these features of TIPPS: H_{2}S + hν → H^{+} + SH^{−} H_{2}S + hν → H^{−} + SH^{+} HCN + hν → H^{+} + CN^{−} Our results on these systems indicate that, as with neutral photodissociation, an initial state of the neutral can be coupled to a wide range of product states, meaning that the spectroscopy of highly excited ions can be studied by TIPPS.